OpenStudy (anonymous):

Find the non-extraneous solutions of the square root of the quantity x plus 6 minus 5 equals quantity x plus 1. x = −2 x = −6 and x = −5 x = −6 and x = −2 x = 2

OpenStudy (anonymous):

@Hero @pooja195 @Nnesha

Nnesha (nnesha):

\[\huge\rm \sqrt{x+6}-5=x+1\]like this ?

OpenStudy (anonymous):

ya

Nnesha (nnesha):

first we need to get rid of `-5` from left side so move -5 to the right side how would you do that ? would you add ,subtract ,multiply or divide ?

OpenStudy (anonymous):

add

Nnesha (nnesha):

yes right add 5 both sides

Nnesha (nnesha):

what did you get ?

OpenStudy (anonymous):

6

Nnesha (nnesha):

yes right so \[\sqrt{x+6}=x+6\] now take square both sides to cancel out the square root \[\huge\rm (\sqrt{x+6})^2= (x+6)^2\]

Nnesha (nnesha):

square root and square cancel each other out you will get x+6 at left side now foil (X+6)^2 (X+6)^2 is same as (x+6)(x+6) multiply

OpenStudy (anonymous):

sorry my internet is being slow

Nnesha (nnesha):

it's okay take ur time :=)

OpenStudy (anonymous):

x^2+36

Nnesha (nnesha):

no that's not right there is a plus sign so you should get 3 terms

Nnesha (nnesha):

 multiply top two terms with the left side write ur answer in the box